# Bayesian Optimization in Language space: An Eval-Efficient AI Self-Improvement Framework

### v1

@@ 4.2 Critic LLM and Self-improving AI algorithm
where $L$ denotes the number of input samples of $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
Given access to a history of past evaluations $H_{t-1}=\{(\Pi_{\tau},s_{\tau})\}_{i=\tau}^{t-1}$ where $s_{\tau}$ is the empirically evaluated score for the prompts $\Pi_{\tau}$, a self-improving AI algorithm is defined as an algorithm $\mathcal{A}$ iterates the following steps at each $t$ to optimize for $J(\Pi)$:
@@
    Evaluation: The system $\Phi(\Pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$. The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},s_{t})\}$.
[…]
Figure 3: A self-improving AI algorithm implemented on the example AI system in Figure 2.
Figure 3 illustrates how the example AI system described in Figure 2 is positioned in the self-improving AI.
5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
We now describe the self-improving AI framework proposed in this paper, TextGrad-Best-of-N Bayesian Optimization (T-BoN BO) (Algorithm 1). Along with the system to optimize $\Phi$ and the scoring rule $r$, it requires a critic model $M_{\text{critic}}$.
Algorithm.
T-BoN BO starts from an initial prompt $\Pi_{0}$. Initially at iteration $t=0$, it gets the initial system $c_{0}=\Phi(\Pi_{0})$ and evaluates it on $\mathcal{D}_{\mathcal{X}}(\Pi_{0})$ to obtain initial score $s_{0}=\mathrm{Eval}(c_{0},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$. The history and the reflection is initialized as $H_{0}=\{(\Pi_{0},c_{0},s_{0})\}$ and $R_{0}=\emptyset$.
We first fix the hyperparameters: the number of candidates per gradient step ($N$), the number of gradient steps per iteration ($G$), and the number of iterations ($T$). Then at each iteration $t$, it first sets $\Pi_{t,0}$ as $\Pi_{t-1}$, the previous iteration’s prompt, and proceeds as follows:
  • 1.
    Best-of-N Textual Gradient steps. (Detailed in Section 5.1) Repeat steps 1) – 4) for $g=1,\ldots,G$:
@@ 5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
      Textual gradient generation: From the current inner prompt $\Pi_{t,g-1}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{(i)}_{t,g}\}_{i=1}^{N}$.
@@
      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{(i)}$ to $\Pi_{t,g-1}$ for all $i$ to get $\{\Pi_{t,g}^{(i)}\}_{i=1}^{N}$.
@@
      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{(i)}:=\Phi(\Pi_{t,g}^{(i)})\}_{i=1}^{N}$.
@@
      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{(i)},c_{t,g}^{(i)})\}_{i=1}^{N}$ and set the best one as $(\Pi_{t,g},c_{t,g})$.
  • 2.
    Evaluation. (Detailed in Section 5.2) Set $c_{t}=c_{t,G}$ and $\Pi_{t}=\Pi_{t,g}$. Evaluate $c_{t}$ to get the evaluation result $s_{t}$, i.e., $s_{t}=\mathrm{Eval}(c_{t},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$. Append $(\Pi_{t},c_{t},s_{t})$ to $H_{t-1}$ to obtain $H_{t}$. If $s_{t}>s_{t-1}$, keep $(\Pi_{t},c_{t})$. Otherwise, set $(\Pi_{t},c_{t})=(\Pi_{t-1},c_{t-1})$ and set $s_{t}=s_{t-1}$. [^6] This ‘gated’ acceptance/rollback was adopted from how textgrad implemented TextGrad.
  • 3.
    Meta-reflection. (Detailed in Section 5.3) Update the reflection $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$ for use at iteration t+1t{+}1.
After $T$ optimization steps, $\{c_{t}\}_{t=1}^{T}$ and $\{s_{t}\}_{t=1}^{T}$ are reported as the algorithm’s optimization outcome.
Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before one evaluation. This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
[…]
Figure 4: Summary of T-BoN BO: $T$ iterations, $G$ gradient steps, and Best-of-$N$ candidates at each gradient step. T-BoN BO deliberately trades inexpensive generation compute for fewer evaluations: each iteration generates $N\times G$ prompt/outcome candidates with no evaluation.
Input: System $\Phi$; initial prompt $\Pi_{0}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$ ; Output: Optimized $\Pi_{T},c_{T},s_{T}$ 1 $c_{0}\leftarrow\Phi(\Pi_{0})$; 2 $s_{0}\leftarrow\text{Eval}(c_{0},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 3 $H_{0}\leftarrow\{(\Pi_{0},c_{0},s_{0})\}$, $R_{0}\leftarrow\emptyset$; 4 for $t=1$ to $T$ do 5 $\Pi_{t,0}\leftarrow\Pi_{t-1}$; 6 for $g=1$ to $G$ do 7    for $i=1$ to $N$ do 8       $\delta^{(i)}_{t,g}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1};R_{t-1},M_{\text{critic}})$ ; 9       $\Pi_{t,g}^{(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1},\delta_{t,g}^{(i)})$ ; 10       $c_{t,g}^{(i)}\leftarrow\Phi(\Pi_{t,g}^{(i)})$ ; 11        12      end for 13    $(\Pi_{t,g},c_{t,g})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{(i)},c_{t,g}^{(i)})\},R_{t-1},M_{\text{critic}})$ 14   end for 15 $c_{t}\leftarrow c_{t,G}$, $\Pi_{t}\leftarrow\Pi_{t,G}$; 16 $s_{t}\leftarrow\text{Eval}(c_{t},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 17 $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$; 18 if $s_{t}<s_{t-1}$ then 19    $\Pi_{t}\leftarrow\Pi_{t-1}$, $c_{t}\leftarrow c_{t-1}$, $s_{t}\leftarrow s_{t-1}$; 20     21   end if 22 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 23  24 end for return $\{c_{t}\}_{t=1}^{T}$ Algorithm 1 T-BoN BO
Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$ 19      end for 20    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 21    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 22    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 23    if $s_{t}^{j}<s_{t-1}^{j}$ then 24       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 25        26      end if 27     28   end for 29 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 30  31 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 2 Parallel T-BoN BO
Figure 5: T-BoN algorithm (Left) and Parallel T-BoN algorithm (Right).
5.1 Best-of-N Textual Gradient steps.
At each gradient step $g=1,\ldots,G$, the algorithm expands the current prompt $\Pi_{t,g-1}$ into a neighborhood of $N$ variants, then contracts back to a single survivor. The process begins with textual gradient generation: the critic model $M_{\text{critic}}$ produces $N$ improvement directions,
$\delta^{(i)}_{t,g}\;\sim\;\nabla_{\text{text}}\!\big(\Pi_{t,g-1};\,R_{t-1},\,M_{\text{critic}}\big),\qquad i=1,\ldots,N,$ (5)
where randomness is introduced by a nonzero sampling temperature and nucleus filtering (top-$p$ sampling) [^7] In the experiments, we used Gemini Flash 2.5 with temperature =0.8=0.8 and top $p=0.9$. , which ensures that the gradients are not identical but represent diverse, locally valid edits.
Each textual gradient is then applied to form a candidate prompt
$\Pi^{(i)}_{t,g}=\operatorname{Apply}\!\big(\Pi_{t,g-1},\,\delta^{(i)}_{t,g}\big)$, (6)
with the rewrite operator configured for stability (low/zero temperature) so the edits remain coherent. The generator model then produces the system outcomes
$c^{(i)}_{t,g}:=\Phi\!\big(\Pi^{(i)}_{t,g}\big)$. (7)
@@
The system outcome candidates are reduced to one winner by the Best-of-N choice operation. For the Best-of-N choice operation, following liu2025pairjudgermperformbestofn, we employ the pairwise tournament that uses critic model $M_{\text{critic}}$. First, the $N$ candidates are permuted to produce randomly chosen tournament matches. Each match compares two outcomes $(c^{(i)}_{t,g},c^{(j)}_{t,g})$ by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 12 in Appendix A). illustrates the prompt example for ad generation tasks. To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove position bias (shi2025judgingjudgessystematicstudy), and break even by random; here we chose $K=4$), producing votes $\{V_{k}(i,j)\}_{k=1}^{K}$, where $V_{k}(i,j)=i$ if critic prefers $i$ over $j$, and vice versa. The winner of the pairwise match is then defined by the majority rule $\operatorname{argmax}_{u\in\{i,j\}}\sum_{k=1}^{K}\mathbf{1}_{\{V_{k}(i,j)=u\}}$. Winners advance round by round until only one candidate remains. The survivor $(\Pi_{t,g},c_{t,g})=(\Pi^{(i^{\star})}_{t,g},c^{(i^{\star})}_{t,g})$ is carried forward for the next gradient step $g+1$.
5.2 Evaluation
After completing $G$ gradient steps, we arrive at the $(\Pi_{t,G},c_{t,G})$, which we redefine as $(\Pi_{t},c_{t})$. We evaluate $(\Pi_{t},c_{t})$: recall from Section 4 that $(\Pi_{t},c_{t})$ is evaluated as
$s_{t}\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c_{t},x_{t}^{(l)}\right)\right),\qquad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi_{t}),$ (8)
where $r\!\left(y\left(c_{t},x_{t}^{(l)}\right)\right)$ is the evaluation of $c_{t}$ for the input $x_{t}^{(l)}$ and $L$ denotes the number of inputs evaluated. After evaluation, the triple $(\Pi_{t},c_{t},s_{t})$ is appended to the history $H_{t-1}$ to form $H_{t}$. Following textgrad, we adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$.
5.3 Meta-reflection.
At the end of each iteration $t$, we compress the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$. $R_{t}$ consists of a small set of actionable rules. In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients. (For the prompt used for meta-reflection, see Figure 13 in Appendix A).
5.4 Parallel T-BoN BO.
Although a local maximum found by gradient steps from a single starting point is often effective in practice, it may under-explore the ample prompt space. Following the practices (wilson2018maximizing; papenmeier2025understanding), we also introduce a parallel variant of T-BoN BO that runs $J$ optimization trajectories simultaneously (Algorithm 2). Here, each of $J$ trajectories starts from a different initial prompt (denoted as $\{\Pi_{0}^{j}\}_{j=1}^{J}$). After independent gradient steps and evaluations at each iteration, evaluations from all trajectories are merged into the shared history, and a single meta-reflection is produced. Different trajectories diversify the search, while the shared reflection ensures that discoveries made by one trajectory are immediately shared with the others. Combined with the careful choice of gradients in T-BoN BO, this parallel process ensures the broader coverage of the search space and faster convergence to high-quality solutions in high-dimensional BO (papenmeier2025understanding) with a slight loss of optimality in the worst case (wilson2018maximizing).
At the end of $T$ optimization steps, the $J$-trajectory parallel T-BoN BO returns $\{(\Pi_{T}^{j},c_{T}^{j},s_{T}^{j})\}_{j=1}^{J}$. For the purpose of reporting, for each $t\leq T$, we select $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$ and report its output $c_{t}^{j_{t}^{\ast}}$ as the step $t$’s outcome. That is, $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$ is reported as the final outcome.
6 Theory: T-BoN BO’s Asymptotic Equivalence to UCB Bayesian Optimization

### v2

@@ 3.2 Self-improving AI algorithm
where $L$ denotes the number of input samples $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
Definition 1 (Self-improving AI algorithm).
Fix a system $\Phi$ and a scoring rule $r$. For each iteration $t\geq 1$, let the history of past evaluations be
$H_{t-1}\;:=\;\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau=0}^{t-1}$ (4)
where $s_{\tau}$ is the empirically evaluated score of $\Pi_{\tau}$, defined as
$s_{t}=\frac{1}{L}\sum_{l=1}^{L}r\bigl(y(\Phi(\Pi_{t}),x_{t}^{(l)})\bigr),\quad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi_{t}).$ (5)
Then, a self-improving AI algorithm is defined as function $\mathcal{A}$ that, at each iteration $t$, maps the history $H_{t-1}$ to a new prompt $\Pi_{t}$, i.e.,
$\Pi_{t}=\mathcal{A}_{t}(H_{t-1})$
The goal of $\mathcal{A}$ is to choose the sequence $(\Pi_{t})_{t\geq 0}$ so as to optimize the objective $J(\Pi)$, equivalently to drive the sequence of empirical scores $(s_{t})_{t\geq 1}$ upward over iterations in expectation.
Operationally, given access to the history $H_{t-1}$, a self-improving AI algorithm $\mathcal{A}$ goes through following three steps each iteration $t$:
@@
    Evaluation: The system outcome $\Phi(\Pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$. The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$.
3.3 Bayesian Optimization (BO)
As discussed in Section 1, we would like to formulate self-improving AI as parallel gradient-based UCB-BO. At its core, BO maximizes an unknown and expensive-to-evaluate objective function under a limited evaluation budget. Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts. BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$. For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
In BO, the following evaluation point is selected by choosing the point that maximizes an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$. Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty). A widely used choice is the Upper Confidence Bound (UCB) acquisition function
$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (6)
where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information). Intuitively, UCB acquisition function gives a positive bonus to unexplored areas in addition to the posterior mean. This allows us to follow the “optimism in the face of uncertainty” principle, which is known to trade off exploration and exploitation effectively. Indeed, UCB-BO is an optimal algorithm for achieving evaluation efficiency, in the sense that no algorithm can achieve better regret guarantees (wang2023regret; whitehouse2023sublinear) (See Web Appendix E for a detailed theoretical discussion of BO.)
In practice, since the maxima of the UCB acquisition function are often not available in a closed form, we typically use parallel gradient-based BO with the UCB acquisition function (wilson2018maximizing). In parallel gradient-based BO, we start from multiple initial points and take gradients from them to climb up the UCB acquisition function to find its local maxima. We then choose the points to evaluate among the local maxima (Figure 3).
In our setting, we do not construct an explicit $\mu(\cdot)$ and $\sigma(\cdot)$ (and their corresponding surrogate model and acquisition function) for the prompt space; instead, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs within the LLM and design our method to emulate parallel gradient-based UCB in this implicit space. Textual gradient (textgrad) is designed to elicit gradients for $\mu(\cdot)$ in the prompt space without explicitly modeling those beliefs within the LLM. However, what we want is not gradients for $\mu(\cdot)$; we want gradients for $A_{\beta}(z)$, the UCB acquisition function.
4 Best-of-N gradient and Theory
As discussed in Section 1, a core challenge in casting self-improving AI as parallel, gradient-based UCB Bayesian optimization in the space of prompts is the absence of an explicit acquisition function (e.g., UCB) and its gradient. As we discussed earlier, textual gradients (textgrad) can guide the prompt toward the highest expected improvement based on immediate feedback, as a textual gradient elicits the gradient of the posterior mean $\nabla\mu(x)$. However, following the posterior mean’s gradient rather than UCB’s is a purely exploitative strategy that often leads to suboptimal local maxima.
In this section, we propose that a gradient direction we call Best-of-N gradient, which generates $N$ stochastic textual gradients and selecting the most promising candidate via a Best-of-$N$ mechanism, can elicit the gradient of the UCB acquisition function within LLM’s mind without explicitly constructing the LLM’s implicit posterior mean and the associated uncertainty.
To operationalize the Best-of-N gradient, we should first formalize the requisite capabilities of the external Large Language Model, which we designate as the Critic LLM.
4.1 Critic LLM.
Let $\Pi$ denote a prompt and $c=\Phi(\Pi)$ denote the corresponding system outcome. We define the history of optimization up to iteration $t$ as $H_{t}=\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau\leq t}$. Then we define the Critic LLM, $M_{\text{critic}}$, as an LLM that can conduct the operations of the following four distinct operators:
  1. 1.
    Pairwise Judge Operator: A function that compares two outcome $c_{1}$ and $c_{2}$, given a reflection $R$ and predicts which outcome will yield a higher true score:
    $\textsc{PairwiseJudge}\big(c_{1},c_{2};R,M_{\text{critic}}\big)\in\{c_{1},c_{2}\}$ (7)
  2. 2.
    Reflection Operator: A function that compresses the entire history into a natural-language summary of rules or guidelines:
    $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ (8)
  3. 3.
    Textual Gradient Operator: A stochastic function that generates a textual gradient $\delta$ (e.g., “Make the lighting warmer”) intended to improve the prompt:
    $\delta\sim\nabla_{\text{text}}(\Pi;R)$ (9)
  4. 4.
    Rewrite Operator: A deterministic function that maps a prompt $\Pi$ and an edit $\delta$ to a new prompt $\Pi^{\prime}$:
    $\Pi^{\prime}=\operatorname{Apply}(\Pi,\delta;M_{\text{critic}})$ (10)
In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), (3) propose prompt improvements based on those guidelines (textgrad; li2025test), and (4) apply prompt improvements to get the new prompt.
4.2 Best-of-N gradient
The notion of a Best-of-$N$ gradient builds on the widely used Best-of-$N$ strategy for test-time alignment, where an LLM trades additional inference-time (test-time) compute for higher-quality outputs by sampling multiple candidate generations and then selecting the one that appears most aligned with a downstream objective (snell2024scaling; beirami2024theoretical). In the standard Best-of-$N$ strategy for prompting, an LLM generates $N$ candidate prompts, a generation model produces $N$ candidate outcomes, and a reward model selects the highest-scoring one. While effective for choosing a single output, our goal is to improve the prompt itself to ensure sustained performance improvement. To achieve this, we introduce the concept of the Best-of-N Gradient. Instead of sampling $N$ candidate prompts directly, we sample $N$ distinct textual gradients (directions of improvement). By applying these $N$ textual edits to the current prompt and selecting the LLM-as-a-judge-predicted most promising candidate solution (Best-of-N), we aim to identify the textual gradient direction for the implicit acquisition function. The procedure operates as follows:
@@ 4.2 Best-of-N gradient
    Textual gradient generation: Using $M_{\text{critic}}$, query the textual-gradient operator $N$ independent times to obtain a set of small edits:
    $\delta^{(i)}\sim\nabla_{\text{text}}(\Pi;R,M_{\text{critic}}),\quad i=1,\dots,N.$ (11)
@@
    Candidate prompt generation: Apply each edit to the current prompt to generate $N$ candidate prompts:
    $\Pi^{(i)}=\operatorname{Apply}(\Pi,\delta^{(i)};M_{\text{critic}}),\quad i=1,\dots,N.$ (12)
@@
    Candidate system outcome generation: Generate the corresponding system outcomes for each candidate prompt:
    $c^{(i)}=\Phi(\Pi^{(i)}),\quad i=1,\dots,N.$ (13)
@@
    Best-of-N selection: Select the best candidate by iteratively applyng PairwiseJudge operation using $M_{\text{critic}}$ as LLM-as-a-judge. We define this selection as Best-of-N operation, which induces the index $i^{\star}$ as the best one among N:
    $i^{\star}=\textsc{Best-of-N}\big(\{(\Pi^{(i)},c^{(i)})\}_{i=1}^{N};R,M_{\text{critic}}\big).$ (14)
    The resulting prompt $\Pi^{(i^{\star})}$ represents the update step taken by the Best-of-N Gradient.
@@
For the Best-of-N selection, we employ the pairwise tournament (liu2025pairjudgermperformbestofn). First, the $N$ candidates are permuted to produce randomly chosen tournament matches. Each match compares two outcomes by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 11 in Appendix A). illustrates the prompt example for ad generation tasks. To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove positional bias (shi2024judging), and ties are broken by random. Winners advance round by round until only the final winner remains. Given the final winner’s index $i^{\star}$, we define $(\Pi^{j}_{t,g},c^{j}_{t,g}):=(\Pi^{j,(i^{\star})}_{t,g},c^{j,(i^{\star})}_{t,g})$ and carry it forward for the next gradient step $g+1$.
4.3 Theory: Best-of-N gradient’s Asymptotic Equivalence to UCB gradient